Suppose a 500 mL beaker is filled to the brim with ethyl alcohol at a temperature of 5 degrees Celsius. Assume the beaker is made of tempered glass and has effectively 0 thermal expansion. What volume in milliters will overflow the beaker when its temperature reaches 21 degrees Celsius? The coefficient of volume expansion got ethyl alcohol is .0011/degrees Celsius. What volume of water in milliliters will overflow under the same conditions? The coefficient of volume expansion for water is .00021/degrees Celsius.
1
Expert's answer
2018-04-09T02:58:10-0400
1) Ethyl alcohol:
ΔV=VβΔT
ΔV=(500)(21-5)(0.0011)=8.8 mL.
2) Water:
ΔV'=Vβ'ΔT
ΔV'=(500)(21-5)(0.00021)=1.7 mL.